What's the way to solve this?
(4a-3)=a+13/a+1

Answer by algebrahouse.com(1659)

(4a - 3) = (a + 13) / (a + 1)
(a + 1)(4a - 3) = a + 13 {multiplied both sides by (a + 1)}
4aČ + a - 3 = a + 13 {used foil method}
4aČ = 16 {subtracted a and added 3 to both sides}
aČ = 4 {divided both sides by 4}
a = 2 or -2 {took the square root of both sides}

For more help from me, visit: www.algebrahouse.com